# Tania Chakraborty, Eylon Caplan, Dan Goldwasser

### v1

@@ Abstract
VIBE: Can a VLM Read the Room?thanks: Preprint, under review.
@@ 1 Introduction
de Saint-Exupéry, The Little Prince
@@
Understanding social dynamics-such as recognizing emotions and their cause-is something humans do intuitively Spunt and Adolphs (2019), yet how we do it remains a challenging question. Even emotion perception, an intuitive ability for most people, poses several levels of complexity Abbas et al. (2024) arising from a variety of factors such as context, cultural variability and channels of perception (language, vision, etc.). Influential psychology and cognitive neuroscience studies Ekman and Friesen (1978); de Gelder (2006) have shown that significant proportions of socially relevant information is contained in non-verbal cues like facial expression and body language. In certain situations Mehrabian and Wiener (1967); Mehrabian (1972), only 7% of meaning is conveyed through words, with 38% coming from tone of voice and 55% through facial expressions! Even within the visual modality alone, emotion recognition is not simply a matter of identifying facial expressions Barrett and Kensinger (2010), as the same facial expression can convey different emotions depending on body posture Aviezer et al. (2012). E.g., even an iconic emotion indicator like a smile, may not always indicate joy Fernández-Dols and Ruiz-Belda (1995). Despite the intuitive ease with which most people perceive emotions, research has shown that this ability relies on complex mechanisms spanning cognition, perception and understanding.
@@
Figure 1: The man smiles sadly, teary eyed. His partner looks at him with sympathy and pats his leg comfortingly. The VLM (InternVL2 26B) correctly identifies the smile, and the woman next to the man, but is not able to interpret the smile correctly.
@@
In recent years there has been increasing interest in AI systems that can achieve human levels of social intelligence. LLMs in particular have shown remarkable promise in such tasks, though they struggle when presented with complex scenarios Li et al. (2024); Liu et al. (2024d); Mou et al. (2024) and are limited to the textual domain. The recent emergence of VLMs Singh et al. (2022); Bai et al. (2023); Bordes et al. (2024); Zhang et al. (2024a) provides an exciting opportunity to study social scenes by connecting textual and visual information and leveraging the strong reasoning capabilities of LLMs. However, while some works have explored the efficacy of using VLMs for social intelligence tasks Yang et al. (2024); Sun et al. (2024); Bhattacharyya and Wang (2025); Liu et al. (2025), this area still remains relatively underexplored.
@@
In this paper, we conduct a thorough investigation of VLMs’ ability to perform social reasoning tasks. We focus on emotion-related inferences drawn from videos that depict a scene’s social dynamics. Our goal is to test whether VLMs can “read the room”, i.e., assess the alignment between their inferences and human-level understanding of the scene. Our assessment is done at two levels of analysis. First, in terms of the visual cues the model identifies as relevant for understanding the social situation, and second in terms of the pragmatic inferences made on top of these cues. Most works have focused on the correctness of the first step, i.e., testing whether models incorrectly hallucinate a visual pattern Li et al. (2023a); Yin et al. (2023); Liu et al. (2024a); Favero et al. (2024); Rawte et al. (2025). In contrast, we focus on the second and as a result expose a previously overlooked limitation in VLMs: even when the VLM correctly identifies a relevant visual cue, it struggles to interpret it correctly. We refer to this as the Visual Social-Pragmatic (VSP) Inference gap. To further elucidate what we mean by this, we use Figure 1 (a frame representing a short video clip) as an example: while a man is clearly smiling, it does not indicate joy, as evident by his slumped posture and teary eyes. Furthermore, the second person in the scene (a pregnant woman in hospital gown) looks at him in sympathy and comforts him. The smile should be interpreted as sad or bittersweet. However, while the VLM correctly identifies the smile, it misinterprets it as evidence for joy.
@@
    Expose the VSP inference gap in VLMs. Propose a task and dataset that isolates the gap and provides a tool for measuring it. [^1] We will release the dataset under MIT license.
@@ 2 Related Work
Multi-modality in Social Science Tasks: Many methods have been proposed for multi-modal emotion prediction that involve parametric training Zheng et al. (2023); Yun et al. (2024); Yang et al. (2023); Huang et al. (2025). Several other approaches combine LLMs with vision models for multi-modal social understanding, e.g. Zhang et al. (2024b); You et al. (2025); Lei et al. (2024) propose prompting strategies to work around the limited reasoning of VLMs. Other works use the reasoning powers of an LLM in conjunction with vision tools Hyun et al. (2024); Kelly et al. (2024); Etesam et al. (2024). Additionally, some datasets have been proposed for exploring models’ capabilities in theory of mind and in emotion interpretation Lin et al. (2025); Chen et al. (2024). The key difference in our work comes from (1) the use of videos for the temporal dimension and (2) the explicit separation that our dataset makes between the problem of hallucination and VSP inference.
Hallucinations in VLMs: VLMs are known to suffer from hallucinations Liu et al. (2024b). Many methods attempt to measure and mitigate hallucinations, either by breaking down outputs Li et al. (2023a); Yin et al. (2023); Petryk et al. (2024), training Ben-Kish et al. (2024); Xie et al. (2024), or decoding algorithms Manevich and Tsarfaty (2024). Visual hallucinations fundamentally differ from VSP in that a description or explanation may be factually correct in isolation (identifying a smile), but provide the wrong pragmatic interpretation, and therefore, incorrect meaning (not realizing it is a sad smile).
Pragmatics: Pragmatics is the study of how context contributes to meaning Morris (1938), and has a rich history in linguistics. Recently, works have sought to understand and improve the pragmatics in LLMs via grounding of various forms Sravanthi et al. (2024); Fried et al. (2023); Mohapatra et al. (2024); White et al. (2024). In the visual domain, VLMs have also been provided with context coming from outside sources, Luo et al. (2024); Li et al. (2023b); Willemsen et al. (2023), while other works address generating contrastive captions using pragmatic inferences Ou et al. (2023); Tsvilodub and Franke (2023). However, in the visual domain, the study of pragmatics in social contexts remains underexplored.
@@ 3.1 Vocabulary
3.2 Diagnostic Task: Emotion Prediction in Multi-Party Conversational Videos
@@
Several works such as Aviezer et al. (2012) and Barrett and Kensinger (2010) show that emotion recognition is not as simple as identifying facial expressions, but rather highly dependent on context. Aviezer et al. (2011) showed that the same facial expression is interpreted differently by people when there are changes in body posture. Based on these works, in order for a VLM to do well at emotion prediction, it would need to be good at identifying relevant Visual Cues in the video, but just as importantly, would need to excel at making the right VSP Inference about the Visual Cues.
@@
3.3 Proposed Task: Visual Social-Pragmatic Inference
@@ 4 Diagnostic Task: Emotion Prediction
The conversations and video clips we used were sourced from the MC-EIU dataset Liu et al. (2024c) which consists of text transcripts of conversations and video clips of the utterances. Since our focus was videos that require VSP Inferences, we used only the videos explicitly labeled with a non neutral emotion. Additionally, it is known that VLMs struggle with long context videos Qu et al. (2025), so we only used videos that were under 4 seconds. The final dataset included 3,536 Chinese video clips (from three different shows) and 5,589 English clips (from two shows). This formed a robust, diverse set of over 9000 videos spanning multiple languages, cultures, and genres.
As illustrated in Figure 2, we had two main diagnostic experimental settings: (1) The VLM acts independently to do the task, and (2) The VLM acts as a perception agent only, and a stronger LLM (GPT-4o-mini) reasons over the visual information. For both settings, we compare performance with the very strong baseline of an LLM (GPT-4o-mini) operating over the text transcript. The motivation behind having the two settings was to probe two different capabilities of the VLM. In setting (1) the goal was to test the VLM’s ability to directly identify the emotion of the speaker. It was not obvious how the VLM would perform compared to the LLM; while the VLM is much smaller in size, it is aided by rich visual data in addition to the text. In contrast, it was reasonable to expect a boost in performance when the VLM had both modalities, compared to either one alone. In setting (2) the goal was to test the ability of the VLM to identify, accurately interpret and communicate what it saw to the LLM (GPT-4o-mini). In this setting, it was reasonable to expect that the additional visual information would boost the performance of the LLM compared to the text only baseline.
@@ 4.1 VLM Independent Setting
We benchmarked the performance of 4 VLMs of sizes varying from 3B to 26B, from 2 different families of models: InternVL2 and Qwen Chen et al. (2025); Qwen et al. (2025). We chose the models based on their performance on vision benchmarks, support for video and the computational cost of running them. As illustrated in Figure 2, we set up the emotion prediction task for each model under 3 different settings: (1) Vision only, (2) Text Only, and (3) Text + Vision. We set the number of frames to be the maximum allowed frames for the smallest-context model (the InternVL2 models), which is 30 frames, and uniformly sampled this many frames for all models to ensure fairness. The 30-frame, 4-second limitations guaranteed at least 7 frames per second, but in most cases was more than 15.
@@ 4.2 VLM-LLM Agent Setting
In this second diagnostic experiment, we used the VLM as a perception agent. We prompted the VLM to describe the speaker’s facial expression and body language at 3 Levels of VSP Inference (see definitions) to get Video Descriptions and fed them into the LLM for classification. In this experiment, we implemented the same three settings as in the previous experiment: V, T, and T + V (Figure 2).
Results: This experiment’s results are shown in Figure 3. The takeaways were: (1) There are strong signals in the vision which the VLM was able to communicate, (2) yet combining the modalities did not lead to significant improvement. Both the visual and textual modalities exhibited strong individual performance, but naively combining them did not yield a clear performance gain—when the VLM operated independently and also when paired with an LLM. Works like Zheng et al. (2023) suggest that directly incorporating vision is difficult because of the noise in the visual domain.
To better understand the results, we broke them down by emotion. As shown in Figure 7 (App. A), the relative performance of each modality varies across emotions. In order to investigate whether the visual information could still be used, we implemented a simple yet effective weighted voting algorithm. The success of the algorithm depended on two key conditions being true: (1) There is complementary information in the Video Descriptions that cannot be recovered from the text transcripts, and (2) The VLM’s performance has a consistent pattern: it struggles with certain emotions and does better at others, even with small sample sizes.
Figure 3: LLM-VLM Agent Experiment Results.
@@
Figure 4: Main steps of the VIBE dataset curation pipeline.
@@
Overall Performance and Takeaways: Results are in Table 2. For both languages for both calibration settings we showed that our simple algorithm led to much better performance than any of the baselines. This validates our assumptions: (1) There is complementary information in video descriptions not recoverable from the text, and (2) The VLM struggles (and succeeds) in a consistent manner across emotions. The next section of the paper investigates why the VLM struggles: hallucinations or VSP Inference mistakes.
Within lang en Within lang zh Cal. zh test en Cal. en test zh
@@ 5.2 VSP Inference Experiments
We benchmarked all VLMs used in the emotion prediction task on our VIBE dataset to illustrate its difficulty. We additionally include the expensive OpenAI GPT-4o mini model as a reference to large, closed VLMs. All models are prompted using CoT, though we include results for standard prompting in App. B. Finally, we had two humans do the VIBE dataset on 100 instances each as comparison. The results can be seen in Table 3.

### v2

# VIBE: Can a VLM Read the Room?

@@ Abstract
VIBE: Can a VLM Read the Room?
@@ 1 Introduction
little-prince, The Little Prince
@@
Understanding social dynamics—such as recognizing emotions and their cause—is something humans do intuitively intuitive-emotions, yet how we do it remains a challenging question. Even emotion perception, an intuitive ability for most people, poses several levels of complexity emotions-are-complex arising from a variety of factors such as context, cultural variability and channels of perception (language, vision, etc.). Influential psychology and cognitive neuroscience studies facs; gelder have shown that significant proportions of socially relevant information is contained in non-verbal cues like facial expression and body language. In certain situations mehrabian1967decoding; mehrabian1972nonverbal, only 7% of meaning is conveyed through words, with 38% coming from tone of voice and 55% through facial expressions! Even within the visual modality alone, emotion recognition is not simply a matter of identifying facial expressions barrett2010context, as the same facial expression can convey different emotions depending on body posture Aviezer2012BodyCues. E.g., even an iconic emotion indicator like a smile, may not always indicate joy smile-is-not-a-smile. Despite the intuitive ease with which most people perceive emotions, research has shown that this ability relies on complex mechanisms spanning cognition, perception and understanding.
@@
Figure 1: In the video, the man smiles sadly, teary eyed. His partner looks at him with sympathy and pats his leg comfortingly. The VLM (InternVL2 26B) correctly identifies the smile, and the woman next to the man, but is not able to interpret the smile correctly.
@@
In recent years there has been increasing interest in AI systems that can achieve human levels of social intelligence. LLMs in particular have shown remarkable promise in such tasks, though they struggle when presented with complex scenarios llm1; llm2; llm3; wu2025identifying and are limited to the textual domain. The recent emergence of VLMs singh2022flava; bai2023qwen; bordes2024introduction; zhang2024vision provides an exciting opportunity to study social scenes by connecting textual and visual information and leveraging the strong reasoning capabilities of LLMs. However, while some works have explored the efficacy of using VLMs for social intelligence tasks yang2024emollm; sun2024face; Bhattacharyya2025EvaluatingVM; liu2025culturevlm, this area still remains relatively underexplored.
@@
In this paper, we conduct a thorough investigation of VLMs’ ability to perform social reasoning tasks. We focus on emotion-related inferences drawn from videos that depict a scene’s social dynamics. Our goal is to test whether VLMs can “read the room”, i.e., assess the alignment between their inferences and human-level understanding of the scene. Our assessment is done at two levels of analysis. First, in terms of the visual cues the model identifies as relevant for understanding the social situation, and second in terms of the pragmatic inferences made on top of these cues. Most works have focused on the correctness of the first step, i.e., testing whether models incorrectly hallucinate a visual pattern pope; woodpecker; liu2024survey; favero2024multi; rawte2025defining. In contrast, we focus on the second and as a result expose a previously overlooked limitation in VLMs: even when the VLM correctly identifies a relevant visual cue, it struggles to interpret it correctly. We refer to this as the Visual Social-Pragmatic (VSP) Inference gap. To further elucidate what we mean by this, we use Figure 1 (a frame representing a short video clip) as an example: while a man is clearly smiling, it does not indicate joy, as evident by his slumped posture and teary eyes. Furthermore, the second person in the scene (a pregnant woman in hospital gown) looks at him in sympathy and comforts him. The smile should be interpreted as sad or bittersweet. However, while the VLM correctly identifies the smile, it misinterprets it as evidence for joy.
@@
    Expose the VSP inference gap in VLMs. Propose a task and dataset that isolates the gap and provides a tool for measuring it.111dataset: https://huggingface.co/datasets/taniarini/vibe
@@
This paper has two main components, presented in Section 4 and Section 5.
  1. 1.
    Section 4 (Diagnostic Emotion Prediction Task): This section builds an understanding of current VLM capabilities on tasks that require social common sense and cognitive reasoning and serves as motivation for why this is an important problem. Section 3.2 gives a high level definition and motivation for this task.
  2. 2.
    Section 5 (Novel VSPI Task): This section describes in detail the construction of the dataset VIBE, and the subsequent results of evaluating several models on it. Section 3.3 provides a high level introduction of the dataset VIBE and the novel VSPI task.
@@ 2 Related Work
Multi-modality in Social Science Tasks: Many methods have been proposed for multi-modal emotion prediction that involve parametric training facial-mmt; telme; yang-etal-2023-self; huang2025emotionqwentraininghybridexperts. Several other approaches combine LLMs with vision models for multi-modal social understanding, e.g. SoV; idealGPT; lei2024largevisionlanguagemodelsemotion propose prompting strategies to work around the limited reasoning of VLMs. Other works use the reasoning powers of an LLM in conjunction with vision tools smile; visionGPT; Etesam_2024. Additionally, some datasets have been proposed for exploring models’ capabilities in theory of mind and in emotion interpretation lin2025feelbreakingboundariesemotional; chen2024theorymindseyereading. The key difference in our work comes from (1) the use of videos for the temporal dimension and (2) the explicit separation that our dataset makes between the problem of hallucination and VSP inference.
Hallucinations in VLMs: VLMs are known to suffer from hallucinations liu2024surveyhallucinationlargevisionlanguage. Many methods attempt to measure and mitigate hallucinations, either by breaking down outputs pope; woodpecker; petryk2024alohanewmeasurehallucination, training ben-kish-etal-2024-mitigating; xie-etal-2024-v, or decoding algorithms manevich-tsarfaty-2024-mitigating. Visual hallucinations fundamentally differ from VSP in that a description or explanation may be factually correct in isolation (identifying a smile), but provide the wrong pragmatic interpretation, and therefore, incorrect meaning (not realizing it is a sad smile).
Pragmatics: Pragmatics is the study of how context contributes to meaning theory-of-signs, and has a rich history in linguistics. Recently, works have sought to understand and improve the pragmatics in LLMs via grounding of various forms pub; llm-prag-survey; mohapatra-etal-2024-evaluating; white-etal-2024-communicate. In the visual domain, VLMs have also been provided with context coming from outside sources, codis; li-etal-2023-multi-modal; willemsen-etal-2023-resolving, while other works address generating contrastive captions using pragmatic inferences contrastive-clip; contrastive-dataset. However, in the visual domain, the study of pragmatics in social contexts remains underexplored.
Multimodal Datasets: There have been excellent datasets that tackle similar problems of social reasoning in multimodal settings. VCR VCR-dataset is one such popular image based common-sense dataset. A closely related class of such datasets are image based emotion prediction datasets such as image-emotion-dataset-1 and image-emotion-dataset-2. VIBE differs from these in 2 main respects: (1) in contrast to images, videos add a layer of complexity due to longer context, subtle temporal effects changing the meaning and (2) VIBE tackles a deeper problem than these tasks by specifically testing the abilities of models to interpret a visual artifact.
@@ 3.1 Vocabulary
3.2 Diagnostic Emotion Prediction Task Overview
@@
Several works such as Aviezer2012BodyCues and barrett2010context show that emotion recognition is not as simple as identifying facial expressions, but rather highly dependent on context. aviezer2011automaticity showed that the same facial expression is interpreted differently by people when there are changes in body posture. Based on these works, in order for a VLM to do well at emotion prediction, it would need to be good at identifying relevant Visual Cues in the video, but just as importantly, would need to excel at making the right VSP Inference about the Visual Cues.
@@
3.3 Novel Visual Social-Pragmatic Inference Task Overview
@@ 4 Diagnostic Task: Emotion Prediction
The conversations and video clips we used were sourced from the MC-EIU dataset mc-eiu which consists of text transcripts of conversations and video clips of the utterances. Since our focus was videos that require VSP Inferences, we used only the videos explicitly labeled with a non neutral emotion. Additionally, it is known that VLMs struggle with long context videos qu2025doesvisionlanguagemodellost, so we only used videos that were under 4 seconds. The final dataset included 3,536 Chinese video clips (from three different shows) and 5,589 English clips (from two shows). This formed a robust, diverse set of over 9000 videos spanning multiple languages, cultures, and genres.
As illustrated in Figure 2, we had two main diagnostic experimental settings: (1) The VLM acts independently to do the task, and (2) The VLM acts as a perception agent only, and a stronger LLM (GPT-4o-mini) reasons over the visual information. For both settings, we compare performance with the very strong baseline of an LLM (GPT-4o-mini) operating over the text transcript. The motivation behind having the two settings was to probe two different capabilities of the VLM. In setting (1) the goal was to test the VLM’s ability to directly identify the emotion of the speaker. It was not obvious how the VLM would perform compared to the stronger LLM; while the LLM has stronger reasoning capabilities, the VLM is aided by rich visual data in addition to the text. In contrast, it was reasonable to expect a boost in performance when the VLM had both modalities, compared to either one alone. In setting (2) the goal was to test the ability of the VLM to identify, accurately interpret and communicate what it saw to the LLM (GPT-4o-mini). In this setting, it was reasonable to expect that the additional visual information would boost the performance of the LLM compared to the text only baseline. For both settings, to give the VLMs the best chance, we sampled the maximum possible frames that our compute resources allowed (30 frames). This decision was based on prior research that showed that performance generally increases as more frames are sampled frame-rate-1; frame-rate-2.
@@ 4.1 VLM Independent Setting
We benchmarked the performance of 4 VLMs of sizes varying from 3B to 26B, from 2 different families of models: InternVL2 and Qwen chen2025expandingperformanceboundariesopensource; qwen2025qwen25technicalreport. We chose the models based on their performance on vision benchmarks, support for video and the computational cost of running them. As illustrated in Figure 2, we set up the emotion prediction task for each model under 3 different settings: (1) Vision only, (2) Text Only, and (3) Text + Vision. We set the number of frames to be the maximum allowed frames for the smallest-context model (the InternVL2 models), which is 30 frames, and uniformly sampled this many frames for all models to ensure fairness. The 30-frame, 4-second limitations guaranteed at least 7 frames per second, but in most cases was more than 15.
@@ 4.2 VLM-LLM Agent Setting
In this second diagnostic experiment, we used the VLM as a perception agent. We prompted the VLM to describe the speaker’s facial expression and body language at 3 Levels of VSP Inference (see definitions) to get Video Descriptions and fed them into the LLM for classification. For some examples of the 3 Levels of VSP Inference, and an experiment confirming the toggling of levels see App. A.1. In this experiment, we implemented the same three settings as in the previous experiment: V, T, and T + V (Figure 2).
Results: This experiment’s results are shown in Figure 3. The takeaways were: (1) There are strong signals in the vision which the VLM was able to communicate, (2) yet combining the modalities did not lead to significant improvement. Both the visual and textual modalities exhibited strong individual performance, but naively combining them did not yield a clear performance gain—when the VLM operated independently and also when paired with an LLM. Works like facial-mmt suggest that directly incorporating vision is difficult because of the noise in the visual domain. To better understand the results, we broke them down by emotion. As shown in Figure 8 (App. A), the relative performance of each modality varies across emotions. We discuss the implications of these results in the next subsection.
Figure 3: Weighted F1-Score for the VLM-LLM Agent experimental setting. The x-axis represents the various types of input to the LLM. Here Text is the text transcript of the conversation, and Visual means the visual cues generated from the VLM when given the conversation video. The Levels correspond to the various VSP levels of the visual cues (see definitions).
4.3 Emotion Prediction Discussion and Implications
From the experiments in sections 4.1 and 4.2, we saw that while VLMs show reasonable strong performance over the visual domain, there was not a clear improvement over the text based performance (Table 1 and Figure 3). This could imply two things: (1) That there is no visual information that would help improve over the text and/or (2) current VLMs do not have the ability to extract and communicate the necessary information accurately. We designed a simple yet effective algorithm to further analyze this: the Weighted Voting Algorithm. The success of the algorithm depended on two key conditions being true: (1) There is complementary information in the Video Descriptions that cannot be recovered from the text transcripts, and (2) The VLM’s performance has a consistent pattern: it struggles with certain emotions and does better at others, even with small sample sizes. We first describe and present results of the algorithm, then discuss its implications:
@@ 4.3 Emotion Prediction Discussion and Implications
Overall Performance and Takeaways: Results are in Table 2. For both languages for both calibration settings we showed that our simple algorithm led to much better performance than any of the baselines. This validates our assumptions: (1) There is complementary information in video descriptions not recoverable from the text, and (2) The VLM struggles (and succeeds) in a consistent manner across emotions. From these results, we can answer the following questions:
  • Is there complementary information in the videos beyond text? $\rightarrow$ Yes
  • Can a VLM extract and communicate this information? $\rightarrow$ Yes (sometimes)
  • Do VLM successes and failures follow systematic patterns? $\rightarrow$ Yes
  • Can these patterns be leveraged to do social and cognitive reasoning tasks better? $\rightarrow$ Yes
Clearly, vision can play an important role in understanding social scenarios, and VLMs are able to (sometimes) give us valuable information that is not recoverable from the text. This encouraging result motivates us to further study exactly what kinds of things does a VLM struggle with? In the next section we show that hallucinations (well known in the literature) are not the only limitation: VLMs also struggle with VSPI. That is, even when they correctly identify visual elements, there are cases when they cannot interpret the correct VSP meaning of what they identified.
Within
lang en
Within
lang zh
Cal. zh
test en
Cal. en
test zh
@@
Figure 4: Main steps of the VIBE dataset curation pipeline.
@@ 5.2 VSP Inference Experiments
We benchmarked all VLMs used in the emotion prediction task on our VIBE dataset to illustrate its difficulty. We additionally include the expensive OpenAI GPT-4o mini model as a reference to large, closed VLMs. All models are prompted using CoT, though we include results for standard prompting in App. B. Finally, we had two humans do the VIBE dataset on 100 instances each as comparison. The results can be seen in Table 3. These results demonstrate that (1) VLMs struggle with VSPI (which comes intuitively to humans), and (2) The question types in VIBE that VLMs struggle with most (fixed and NT) are consistently hard for all the models we evaluated (but not for humans!). This result underscores the importance of targeting VSPI for VLMs in order to make them socially and emotionally intelligent. In section 5.3 we dive into a deeper analysis of the kinds of mistakes made on VIBE and connect the results back to the emotion prediction task from Section 4